OpenStudy (anonymous):

solve for x 1/x ^2 + 1/x = 6

OpenStudy (anonymous):

x=1/2 or -1/3 get a common denominator, and you eventually get to 0=6x^2-x-1 that factors to 0=(3x+1)(2x-1)

OpenStudy (anonymous):

ok let me try brb g\cause I got a couple diff answers

OpenStudy (anonymous):

is the common denominator not x ^2 so I have 1/x^2 + x/x^2 =6x?

OpenStudy (anonymous):

you don't multiply the 6 by x

OpenStudy (anonymous):

it's just 1/x^2 + x/x^2 =6

OpenStudy (anonymous):

ooo right so then 1/x^2 + x/x^2 =6x^2

OpenStudy (anonymous):

x^2 multiplied on everything

OpenStudy (anonymous):

so then I get x+1/x^2 = 6x^2 and I feel like I need to multiply x^2 on both sides and then bring everything on one side but it starts to look ugly am I headed in the wrong direction?

OpenStudy (anonymous):

you don't need to multiply EVERYTHING by x^2, just 1/x. 1/x = x/x^2, just a different way of writing it. the equation still stays balanced.

OpenStudy (anonymous):

I thought when u have this its just like addition with fractions and u need to find a common denominator. So what should my equation look like